# An Extensive Study on D2C: Overfitting Remediation in Deep Learning Using a Decentralized Approach

### v1

@@ VI-C Accuracy, F1 Score, AUC-ROC, and Log Loss
Model Settings Accuracy F1-Score Log Loss AUC-ROC
No Subset 0.9243 0.9243 0.2611 0.98764
N=2, E=1 0.9291 0.9291 0.2556 0.98873
N=3, E=1 0.9288 0.9288 0.2520 0.98872
N=3, E=2 0.9275 0.9275 0.2567 0.98852
N=5, E=1 0.9293 0.9292 0.2412 0.98881
N=5, E=2 0.9295 0.9296 0.2451 0.98875
N=7, E=1 0.9271 0.9271 0.2353 0.98916
@@
We can observe from Table VII that, once again, the highest accuracy and F1-score were achieved with our proposed D2C method. The number of subsets, $N$, was $5$ in this case for the best model. The number of epochs before each round of central averaging, $E$, was $2$. The D2C method showed a significant improvement with an accuracy of $92.95\%$ over the traditional approach with no subsets, which yielded an accuracy of $92.43\%$. Using the D2C method decreased the error by more than $0.5\%$ in this case. Also, as we can see from Table VII, the trend of a decreasing log loss as we increase the number of subsets is visible here too. The best method in terms of AUC-ROC also happens to be the model with the highest number of subsets (7). It also seems that the log loss increases a bit when we increase the number of epochs, $E$.
@@
The effect of dividing the training set into multiple subsets is the most apparent in this dataset compared to all the previous ones. AG News, the largest dataset in our study, presents an ideal case for D2C, as each class contains 30,000 samples. This abundance of data per class allows for substantial division without sacrificing subset size, which maintains the model’s training stability while leveraging the regularizing effect of partitioning and averaging. Consequently, every variation of D2C (across all values of $N$ and $E$) demonstrated accuracy gains over the traditional approach, underscoring D2C’s suitability for mitigating overfitting in text classification tasks.
@@
In summary, D2C excels with large datasets like AG News, where ample samples per class allow the model to benefit from both training on robust subsets and reducing overfitting through partitioning and averaging. This validates D2C’s potential for significant impact in Big Data applications, where traditional models often struggle to fully exploit available data.
VI-D Augmentation in D2C
@@
Our experiments show that dividing training data into multiple subsets and applying central averaging results in less overfitting and an improvement in performance metrics like accuracy, F1 score, etc. However, we also saw that the training subsets need to be sufficiently large so that each of the subsets can provide enough training diversity to the edge models, otherwise, the performance suffers a drop. When the dataset is large, the proposed method works very well, but when the dataset is smaller, this may cause issues in achieving optimum performance.
@@
One way to counter these issues is data augmentation. This is a widely used technique to expand a dataset and prevent overfitting. To minimize the effect of outliers and noise and to prevent data leakage among the subsets, augmentation must take place after dividing the dataset into multiple subsets. The augmentation is then applied to each of the subsets. That way, the augmented samples from one subset won’t get mixed with another subset. This approach ensures the separation of data in each of the subsets as well as the expansion of training data, providing enough samples for all the subsets, thus improving the overall performance. We tested this augmentation approach on the CIFAR-10 dataset for $3$, $5$ $7$ subsets. The augmentation methods we used were rotation, horizontal flip, and shifting along both axes. Let us take a look at the accuracies we achieved in each case and compare them with the accuracies we got before without applying augmentation.
@@
Figure 15: Test accuracies for different numbers of subsets, N, using the D2C method on the CIFAR-10 dataset with and without Augmentation.
@@
As we can see from Figure 15, the accuracy improved significantly in each case after we applied data augmentation to the subsets of the training data. The accuracy increased from $86.13\%$ to $88.73\%$ for $3$ subsets. In the case of $5$ subsets, the improvement was even more, reaching $88.12\%$ from just $85.52\%$. In the case of $7$ subsets too, the accuracy increased from $85.39\%$ to $87.58\%$. This shows that data augmentation techniques can be successfully applied when using the D2C method, and they can effectively treat the problems that arise when the size of the subsets becomes too small.
VII Limitations
@@
We have shown that our proposed D2C method treats overfitting and shows an improvement over the traditional approach in multiple domains and datasets. However, all these come at a cost. Dividing the training set into multiple subsets and training them all separately means repeating the whole training process $N$ times ($N$ = number of subsets), albeit with less data. This results in an increased training time. Let us take a look at the following Figure 16 for MNIST.
@@
Figure 16: Time elapsed per epoch for the traditional approach and different numbers of subsets, N, using the D2C method on the MNIST dataset.
@@
The training was done using a V100 GPU in each of these cases for fair comparisons. As we can see from Figure 16, the time needed to train the models is the least in the traditional approach, as expected. As we increase the number of subsets, $N$, the time needed to complete each epoch of training increases. However, it is important to note that the increase in training time is not abrupt. It doesn’t increase proportionally with $N$. For example, the time taken to train for one epoch was $12.5$ seconds in the case of the traditional implementation. When the number of subsets, $N$, was $11$, it took $40$ seconds to train for one global epoch. The training time increased significantly, but it was nowhere near $11$ times the time taken in the traditional approach. So, even though the training time increases, it doesn’t shoot up abruptly. It is important because such an abrupt increase in training time could mean a big issue when it comes to applying the Divide2Conquer method.
@@
Lastly, it’s important to note that even if the training time is more when using the D2C method, it doesn’t make any difference at all in the post-training deployment stage, as the edge models (like the one used in the traditional approach) and the central model have the same number of weights, architecture, and computational complexity. So, clearly, even if the Training time increases while using the Divide2Conquer method, the Inference time remains exactly the same as the traditional approach.
VIII Conclusion
@@
Our study demonstrates that dividing training data into multiple subsets and averaging the weights iteratively can treat overfitting and produce significant performance gains through stability and better generalization. In this paper, we applied our proposed Divide2Conquer method and tested it on multiple datasets from different domains. This method outperformed the traditional approach in all the datasets in terms of key metrics and showed significantly better resistance against overfitting. The D2C method also showed a consistent ability to mitigate the increasing trend of validation loss, and to produce smoother decision boundaries. These results prove our hypothesis of this method being able to help generalize better and resist overfitting to be true. Moreover, the Divide2Conquer method brings even more improvement when the dataset is larger, outlining its promise in Big Data scenarios, especially since D2C method can be applied on top of other generalization techniques. Thus, the proposed D2C method opens the door for future research and experiments using this in several domains.
Acknowledgments
@@
Appendix A Related Proofs and Derivations
Derivation of $Var(X)=\mathop{{}\mathbb{E}}(X^{2})-(\mathop{{}\mathbb{E}}(X))^{2}$ : Let the random variable $X$ denote the estimate of $x$. The variance of estimating this random variable is defined as the average deviation of $X$ from the expected value of our estimate, $\mathop{{}\mathbb{E}}(X)$, where the deviation is squared for symmetry of difference. So, the variance can be formulated as:
@@ Appendix A Related Proofs and Derivations
$=\mathop{{}\mathbb{E}}(X^{2}+(\mathop{{}\mathbb{E}}(X))^{2}-2X\mathop{{}\mathbb{E}}(X))$
=$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\mathop{{}\mathbb{E}}(X^{2})+(\mathop{{}\mathbb{E}}(X))^{2}-2\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(\mathop{{}\mathbb{E}}(X))$
$=\mathop{{}\mathbb{E}}(X^{2})-(\mathop{{}\mathbb{E}}(X))^{2}$
Here, $(a)$ is due to the linear nature of the expectation operator, and $\mathop{{}\mathbb{E}}(\mathop{{}\mathbb{E}}(X))=\mathop{{}\mathbb{E}}(X)$, because $\mathop{{}\mathbb{E}}(X)$ is not a random variable.
Derivation of $CoV(XY)=\mathop{{}\mathbb{E}}(XY)-\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(Y)$ : Let $X$ and $Y$ be estimations of two random variables $x$ and $y$ respectively. The covariance of $X$ and $Y$ is defined as the expected value (or mean) of the product of their deviations from their individual expected values[35]. Hence,
$Cov(X,Y)$ $=\mathop{{}\mathbb{E}}((X-\mathop{{}\mathbb{E}}(X)).(Y-\mathop{{}\mathbb{E}}(Y)))$
$=\mathop{{}\mathbb{E}}(XY-X\mathop{{}\mathbb{E}}(Y)-Y\mathop{{}\mathbb{E}}(X)+\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(Y))$
=$\stackrel{{\scriptstyle\text{(b)}}}{{=}}\mathop{{}\mathbb{E}}(XY)-\mathop{{}\mathbb{E}}(X\mathop{{}\mathbb{E}}(Y))$
$\qquad-\mathop{{}\mathbb{E}}(Y\mathop{{}\mathbb{E}}(X))+\mathop{{}\mathbb{E}}(\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(Y))$
=$\stackrel{{\scriptstyle\text{(c)}}}{{=}}\mathop{{}\mathbb{E}}(XY)-\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(Y)$
$\qquad-\mathop{{}\mathbb{E}}(Y)\mathop{{}\mathbb{E}}(X)+\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(Y)$
$=\mathop{{}\mathbb{E}}(XY)-\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(Y)$
@@
[…]
=$\stackrel{{\scriptstyle\text{(d)}}}{{=}}\int yf(y)\mathop{{}\mathbb{E}}(x)dy$
[…]
$=\mathop{{}\mathbb{E}}(X)E(Y)$
@@
Let us consider the random variables $X$ and $Y$ and arbitrary constants $a,b$ Using LABEL:proof1 we can write,
$Var(aX+bY)$ $=\mathop{{}\mathbb{E}}((aX+bY)^{2})-(\mathop{{}\mathbb{E}}(aX+bY))^{2}$
$=(a^{2}\mathop{{}\mathbb{E}}(X^{2})+b^{2}\mathop{{}\mathbb{E}}(Y^{2})+2ab\mathop{{}\mathbb{E}}(XY))$
$\qquad-(a\mathop{{}\mathbb{E}}(x)+b\mathop{{}\mathbb{E}}(Y))^{2}$
$=(a^{2}\mathop{{}\mathbb{E}}(X^{2})+b^{2}\mathop{{}\mathbb{E}}(Y^{2})+2ab\mathop{{}\mathbb{E}}(XY))$
$\qquad-a^{2}(\mathop{{}\mathbb{E}}(X))^{2}-b^{2}(\mathop{{}\mathbb{E}}(Y))^{2}$
$\qquad-2ab\mathop{{}\mathbb{E}}(X)\mathop{{}\mathbb{E}}(Y)$
=$\stackrel{{\scriptstyle\text{(e)}}}{{=}}a^{2}Var(X)+b^{2}Var(Y)+2abCov(X,Y)$

### v2

@@ 5.3. Accuracy, F1 Score, AUC-ROC, Cohen’s Cappa, Log Loss and MCC
Model Settings Accuracy F1-Score ROC-AUC Cohen Kappa Log Loss Matthews CC
No Subsets 0.9243 0.9243 0.98764 0.8991 0.2611 0.8992
N=2, E=1 0.9300 0.9300 0.98856 0.9067 0.2519 0.9067
N=3, E=1 0.9312 0.9312 0.98888 0.9082 0.2378 0.9084
N=3, E=2 0.9278 0.9277 0.98879 0.9037 0.2411 0.9037
N=5, E=1 0.9274 0.9273 0.98905 0.9032 0.2416 0.9032
N=5, E=2 0.9287 0.9287 0.98830 0.9049 0.2656 0.9049
N=7, E=1 0.9288 0.9288 0.98836 0.9051 0.2443 0.9052
@@
We can observe from Table 7 that, once again, the highest accuracy and F1-score were achieved with our proposed DUA-D2C method. The number of subsets, $N$, was $3$ in this case for the best model. The number of epochs before each round of central averaging, $E$, was $1$. The DUA-D2C method showed a significant improvement with an accuracy of $93.12\%$ over the traditional approach with no subsets, which yielded an accuracy of $92.43\%$. The best method in terms of AUC-ROC happens to be the model with even more subsets (5). It also seems that the log loss increases a bit when we increase the number of epochs, $E$.
@@
The effect of using DUA-D2C is the most apparent in this dataset compared to all the previous ones. AG News, the largest dataset in our study, presents an ideal case for DUA-D2C, as each class contains 30,000 samples. This abundance of data per class allows for substantial division without sacrificing subset size, which maintains the model’s training stability while leveraging the regularizing effect of partitioning and averaging. Consequently, every variation of DUA-D2C (across all values of $N$ and $E$) demonstrated accuracy gains over the traditional approach, underscoring DUA-D2C’s suitability for mitigating overfitting in text classification tasks.
@@
In summary, DUA-D2C excels with large datasets like AG News, where ample samples per class allow the model to benefit from both training on robust subsets and reducing overfitting through partitioning and averaging. This validates DUA-D2C’s potential for significant impact in Big Data applications, where traditional models often struggle to fully exploit available data.
Table 8. Comparison of Best Performance: Traditional vs. D2C vs. DUA-D2C across Datasets for Key Metrics.
Dataset Metric Traditional D2C (Best) DUA-D2C (Best)
CIFAR-10 Accuracy 0.8604 0.8613 0.8652
F1-Score 0.8601 0.8611 0.8648
ROC-AUC 0.98731 0.98704 0.98761
Log Loss 0.7198 0.5733 0.5876
Fashion MNIST Accuracy 0.9312 0.9338 0.9344
F1-Score 0.9313 0.9337 0.9340
ROC-AUC 0.99546 0.99600 0.99600
Log Loss 0.3492 0.2341 0.2314
MNIST Accuracy 0.9952 0.9959 0.9963
F1-Score 0.9952 0.9959 0.9963
ROC-AUC 0.99995 0.99999 0.99999
Log Loss 0.0258 0.0161 0.0148
FER2013 Accuracy 0.6498 0.6567 0.6537
F1-Score 0.6420 0.6465 0.6457
ROC-AUC 0.89934 0.90320 0.90528
Log Loss 2.0112 1.3941 1.1934
Audio_Combined Accuracy 0.6953 0.6942 0.7010
F1-Score 0.6870 0.6909 0.6973
ROC-AUC 0.96570 0.96553 0.96560
Log Loss 1.3419 1.0033 1.0571
AG News Accuracy 0.9243 0.9295 0.9312
F1-Score 0.9243 0.9296 0.9312
ROC-AUC 0.98764 0.98916 0.98905
Log Loss 0.2611 0.2353 0.2378
  • Values in D2C (Best) and DUA-D2C (Best) columns represent the optimal performance across all tested N (Number of Subsets) and E (Local Epochs) configurations for that method. Bold indicates the best overall performance among the three approaches (Traditional, D2C, DUA-D2C) for each metric and dataset. For Log Loss, lower values are better.
The results presented in Table 8 demonstrates the advantages of the ”divide and conquer” strategy. The D2C method consistently outperforms the traditional monolithic training approach across various datasets and key metrics, underscoring the inherent benefits of partitioned training and parameter averaging for improved generalization.
Building upon this foundation, the DUA-D2C method further refines and perfects this technique. Through its dynamic, uncertainty-aware aggregation, DUA-D2C consistently achieves superior performance compared to both the traditional method and the standard D2C. While the D2C method itself provides substantial gains, the intelligent weighting scheme of DUA-D2C offers an additional performance enhancement. This positions DUA-D2C as the more advanced and robust implementation of our proposed framework, consistently delivering strong, and in many cases, the best, generalization capabilities.
5.4. Augmentation in DUA-D2C
@@
Our experiments show that applying DUA-D2C results in less overfitting and an improvement in performance metrics like accuracy, F1 score, etc. However, we also saw that the training subsets need to be sufficiently large so that each of the subsets can provide enough training diversity to the edge models, otherwise, the performance suffers a drop. When the dataset is large, the proposed method works very well, but when the dataset is smaller, this may cause issues in achieving optimum performance.
@@
One way to counter these issues is data augmentation. This is a widely used technique to expand a dataset and prevent overfitting. To minimize the effect of outliers and noise and to prevent data leakage among the subsets, augmentation must take place after dividing the dataset into multiple subsets. The augmentation is then applied to each of the subsets. That way, the augmented samples from one subset won’t get mixed with another subset. This approach ensures the separation of data in each of the subsets as well as the expansion of training data, providing enough samples for all the subsets, thus improving the overall performance. We tested this augmentation approach on the CIFAR-10 dataset for $3$, 55 & $7$ subsets. The augmentation methods we used were simple rotation, horizontal flip, and shifting along both axes. Let us take a look at the accuracies we achieved in each case and compare them with the accuracies we got before without applying augmentation.
@@
Figure 14. Test accuracies for different numbers of subsets, N, using the DUA-D2C method on the CIFAR-10 dataset with and without Augmentation.
@@
As we can see from Figure 14, the accuracy improved significantly in each case after we applied data augmentation to the subsets of the training data. The accuracy increased from $86.52\%$ to $88.74\%$ for $3$ subsets. In the case of $5$ subsets, the accuracy reached $88.12\%$ from just $86.14\%$. In the case of $7$ subsets too, the accuracy increased from $85.56\%$ to $87.68\%$. This shows that data augmentation techniques can be successfully applied when using the DUA-D2C method, and they can effectively treat the problems that arise when the size of the subsets becomes too small.
6. Limitations
@@
We have shown that our proposed DUA-D2C method treats overfitting and shows an improvement over the traditional approach in multiple domains and datasets. However, all these come at a cost. Dividing the training set into multiple subsets and training each subset separately means repeating the entire training process $N$ times ($N$ = number of subsets), albeit with less data. This results in an increased training time. Let us take a look at the following Figure 15 for MNIST.
@@
Figure 15. Time elapsed per epoch for the traditional approach and different numbers of subsets, N, using the DUA-D2C method on the MNIST dataset.
@@
The training was done using a V100 GPU in each of these cases for fair comparisons. As we can see from Figure 15, the time needed to train the models is the least in the traditional approach, as expected. As we increase the number of subsets, $N$, the time needed to complete each epoch of training increases. However, it is important to note that the increase in training time is not abrupt. It doesn’t increase proportionally with $N$. For example, the time taken to train for one epoch was $12.5$ seconds in the case of the traditional implementation. When the number of subsets, $N$, was $11$, it took $43$ seconds to train for one global epoch. The training time increased significantly, but it was nowhere near $11$ times the time taken in the traditional approach. So, even though the training time increases, it doesn’t shoot up abruptly. It is important because such an abrupt increase in training time could mean a big issue when it comes to applying the Divide2Conquer method. However, the inherent parallelism of the D2C/DUA-D2C framework lends itself well to distributed computing environments; deploying edge model training across multiple GPUs or compute nodes could substantially mitigate this increase in training time.
@@
Lastly, it’s important to note that even if the training time is more when using the DUA-D2C method, it doesn’t make any difference at all in the post-training deployment stage, as the edge models (like the one used in the traditional approach) and the central model have the same number of weights, architecture, and computational complexity. So, clearly, even if the Training time increases while using the Divide2Conquer technique, the Inference time remains exactly the same as the traditional approach.
7. Related works
Several authors have addressed the issue of overfitting while performing various tasks. M. Cogswell et al. proposed a regularizer called DeCov, which helps reduce overfitting in deep neural networks by Decorrelating Representations (cogswell2015reducing, 5).
Dropout (srivastava2014dropout, 27) was proposed by Srivastava et al. back in 2014, and since then, this technique has been extensively used in very complex neural network architectures successfully. It was indeed an outstanding contribution, specifically for deep learning-based models. Batch Normalization (ioffe2015batch, 10) proposed by Ioffe and Sergei primarily focuses on better convergence and somewhat contributes to reducing overfitting. J. Kolluri and V.K Kotte came up with L¼ regularization to solve the problems faced by L1 and L2 regularization techniques (kolluri2020reducing, 12).
Ensembles are also often used to improve generalization. The ensemble-based ELVD model (shanmugavel2023novel, 25) managed to outperform the traditional VGGNet and DropoutNet models in terms of reducing overfitting. Zehong Zeng et al. came up with an ensemble framework that incorporates several techniques to prevent overfitting (zeng2023ensemble, 32). Min-Gu Kim et al. also proposed parallel ensemble networks to reduce overfitting in ECG data and prevent the degradation of generalization performance as the training progresses (kim2020ensemble, 11).
We implemented a method based on federated optimization preliminarily for facial expression recognition using FedNet(siddiqui2022fednet, 26). This model achieves excellent results in terms of generalization on the CK+ dataset. The K-Fold-Cross-Validation method for evaluation proposed by Korjus K. et al. (korjus2016efficient, 14) has some similarities to our proposed method in the sense that it also creates subsets by partitioning the dataset and makes use of the whole training data. However, this method is different from ours since the erroneous samples still make it to the training phase (K-1/K)-th of the time in K-Fold CV. The proposed DUA-D2C method partitions data, and the whole training process is partitioned too, which means none of the subsets is fed with a particular outlier while training bar one, minimizing the impact of these erroneous samples.
Generalization has also been studied in advanced architectures like GCNs. Complexity is preserved while introducing measures such as norm regularization and feature projection, as shown by Giovanni Donghi et al. (randgcn2024, 6).
Recent theoretical advances have deepened our understanding of generalization in deep learning. Masegosa and Ortega (masegosa2025pacchernoff, 20) proposed PAC-Chernoff bounds to characterize generalization in the interpolation regime, revealing how smoothness arising from regularization, data augmentation, and architectural invariance can lead to more reliable learning. We find this framework particularly compelling, as it provides a rigorous lens through which to interpret the benefits of our uncertainty-aware aggregation in DUA-D2C. By dynamically favoring models with smoother and more confident predictions, our method echoes the principles outlined in their theoretical analysis.
Similarly, Lyu et al. (lyu2024generalizationgap, 19) investigated the generalization gap in visual reinforcement learning, introducing a representation distance metric to quantify robustness across environments. Although their work focuses on reinforcement learning, it underscores the importance of representation stability, a concept that resonates with our Divide2Conquer philosophy. By training models on diverse data partitions and aggregating them based on validation performance, DUA-D2C implicitly promotes representational diversity while mitigating overfitting.
We deeply appreciate these contributions for their clarity and rigor, and we view DUA-D2C as a practical extension of these theoretical insights into the supervised learning domain.
Overfitting remains a prevalent issue in supervised machine learning despite methods like Early Stopping, Network Reduction, Training Set Expansion, Regularization, and Dropout being effective (ying2019overview, 30). Through DUA-D2C, we build an approach that focuses on achieving even better generalization and can be implemented on top of other overfitting-reducing techniques.
8. Conclusion
@@
Our study lays the theoretical groundwork and demonstrates experimentally that our proposed Dynamic Uncertainty-Aware Divide2Conquer (DUA-D2C) method effectively treats overfitting and produces significant performance gains through improved stability and better generalization. In this paper, we applied our proposed method and tested it on multiple datasets from different domains. This advanced method consistently outperformed both the traditional approach and a simpler D2C averaging scheme in terms of key metrics and showed significantly better resistance against overfitting. The DUA-D2C method also showed a consistent ability to mitigate the increasing trend of validation loss and to produce smoother decision boundaries. These results prove our hypothesis of this refined method being able to help generalize better and resist overfitting to be true. Moreover, the DUA-D2C method often brings even more improvement when the dataset is larger, outlining its promise in Big Data scenarios, especially since the DUA-D2C framework can be readily applied on top of other generalization techniques. Thus, the proposed DUA-D2C method opens the door for future research and experiments, furthering the exploration of intelligent distributed learning in several domains.
Acknowledgements.
@@
Availability of Data and Codes
The datasets analyzed during the study are publicly available. The source codes for this study are available on GitHub at https://github.com/Saiful185/DUA-D2C.
Appendix A Appendix: Related Proofs and Derivations
Derivation of $Var(X)=\operatorname{\mathbb{E}}(X^{2})-(\operatorname{\mathbb{E}}(X))^{2}$ : Let the random variable $X$ denote the estimate of $x$. The variance of estimating this random variable is defined as the average deviation of $X$ from the expected value of our estimate, $\operatorname{\mathbb{E}}(X)$, where the deviation is squared for symmetry of difference. So, the variance can be formulated as:
@@ Appendix A Appendix: Related Proofs and Derivations
$=\operatorname{\mathbb{E}}(X^{2}+(\operatorname{\mathbb{E}}(X))^{2}-2X\operatorname{\mathbb{E}}(X))$
$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\operatorname{\mathbb{E}}(X^{2})+(\operatorname{\mathbb{E}}(X))^{2}-2\operatorname{\mathbb{E}}(X)\operatorname{\mathbb{E}}(\operatorname{\mathbb{E}}(X))$
$=\operatorname{\mathbb{E}}(X^{2})-(\operatorname{\mathbb{E}}(X))^{2}$
Here, $(a)$ is due to the linear nature of the expectation operator, and $\operatorname{\mathbb{E}}(\operatorname{\mathbb{E}}(X))=\operatorname{\mathbb{E}}(X)$, because $\operatorname{\mathbb{E}}(X)$ is not a random variable.
Derivation of $CoV(XY)=\operatorname{\mathbb{E}}(XY)-\operatorname{\mathbb{E}}(X)\operatorname{\mathbb{E}}(Y)$ : Let $X$ and $Y$ be estimations of two random variables $x$ and $y$ respectively. The covariance of $X$ and $Y$ is defined as the expected value (or mean) of the product of their deviations from their individual expected values(park2018fundamentals, 22). Hence,
$Cov(X,Y)$ $=\operatorname{\mathbb{E}}((X-\operatorname{\mathbb{E}}(X)).(Y-\operatorname{\mathbb{E}}(Y)))$
$=\operatorname{\mathbb{E}}(XY-X\operatorname{\mathbb{E}}(Y)-Y\operatorname{\mathbb{E}}(X)+\operatorname{\mathbb{E}}(X)\operatorname{\mathbb{E}}(Y))$
$\stackrel{{\scriptstyle\text{(b)}}}{{=}}\operatorname{\mathbb{E}}(XY)-\operatorname{\mathbb{E}}(X\operatorname{\mathbb{E}}(Y))$
$\qquad-\operatorname{\mathbb{E}}(Y\operatorname{\mathbb{E}}(X))+\operatorname{\mathbb{E}}(\operatorname{\mathbb{E}}(X)\operatorname{\mathbb{E}}(Y))$
$\stackrel{{\scriptstyle\text{(c)}}}{{=}}\operatorname{\mathbb{E}}(XY)-\operatorname{\mathbb{E}}(X)\operatorname{\mathbb{E}}(Y)$
$\qquad-\operatorname{\mathbb{E}}(Y)\operatorname{\mathbb{E}}(X)+\operatorname{\mathbb{E}}(X)\operatorname{\mathbb{E}}(Y)$
$=\operatorname{\mathbb{E}}(XY)-\operatorname{\mathbb{E}}(X)\operatorname{\mathbb{E}}(Y)$
@@
$=\int yf(y)\int xf(x)dxdy$
$\stackrel{{\scriptstyle\text{(d)}}}{{=}}\int yf(y)\operatorname{\mathbb{E}}(x)dy$
$=\operatorname{\mathbb{E}}(X)\int yf(y)dy$
$=\operatorname{\mathbb{E}}(X)E(Y)$
@@
Let us consider the random variables $X$ and $Y$ and arbitrary constants $a,b$ . Using LABEL:proof1 we can write,
$Var(aX+bY)$ $=\operatorname{\mathbb{E}}((aX+bY)^{2})-(\operatorname{\mathbb{E}}(aX+bY))^{2}$
$=(a^{2}\operatorname{\mathbb{E}}(X^{2})+b^{2}\operatorname{\mathbb{E}}(Y^{2})+2ab\operatorname{\mathbb{E}}(XY))$
$\qquad-(a\operatorname{\mathbb{E}}(x)+b\operatorname{\mathbb{E}}(Y))^{2}$
$=(a^{2}\operatorname{\mathbb{E}}(X^{2})+b^{2}\operatorname{\mathbb{E}}(Y^{2})+2ab\operatorname{\mathbb{E}}(XY))$
$\qquad-a^{2}(\operatorname{\mathbb{E}}(X))^{2}-b^{2}(\operatorname{\mathbb{E}}(Y))^{2}$
$\qquad-2ab\operatorname{\mathbb{E}}(X)\operatorname{\mathbb{E}}(Y)$
$\stackrel{{\scriptstyle\text{(e)}}}{{=}}a^{2}Var(X)+b^{2}Var(Y)+2abCov(X,Y)$